Circular No. 8970

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

R/2006 S 5 AND R/2007 S 1 C. C. Porco, CICLOPS, Space Science Institute, Boulder, on behalf of

the Cassini Imaging Team, reports the detection of a ring arc (designated R/2007 S 1) associated with the small satellite Saturn XLIX (Anthe; semimajor axis 197655 km; cf. IAUCs 8857, 8873), and the apparent confirmation of a previously reported ring arc R/2006 S 5 (IAUC 8773) associated with the satellite Saturn XXXII (Methone; semi-major axis 194230 km; IAUCs 8389, 8471). The ring arc in the orbit of Anthe was detected in a series of five 15-s-exposure recovery images of Anthe taken with the Narrow Angle Camera of the Cassini Imaging Science Subsystem on 2007 Oct. 29 at a phase angle of 23^{o}. The satellite appears to be embedded in the arc but is drifting slowly with respect to it. The Anthe arc extends over ∼ 20^{o} in longitude and appears to be associated with material trapped in the 11:10 co-rotation eccentricity resonance with Saturn I (Mimas). Furthermore, additional detections of the Anthe arc between 2008 February and July show that the satellite’s motion with respect to the arc is consistent with Anthe’s ±7^{o} libration in the same 11:10 resonance, suggesting that Anthe is the likely source of the arc material, and also that it may be creating structure within the arc. The earliest detection of the Anthe arc was on 2007 June 15. Serendipitously, the discovery image of the Anthe arc also revealed that the satellite Methone is embedded in an arc of material with a longitudinal extent of ∼ 10^{o}, which may be associated with the 15:14 corotation eccentricity resonance with Mimas. Additional detections of the arc were made in 2008 Feb.–Apr.; the earliest detection was on 2007 May 31. The observed Methone arc could help to explain data from the LEMMS sensor of the Cassini Magnetospheric Imaging Instrument that led to the detection of R/2006 S 5 (IAUC 8773).

COMET P/2008 Q4 (LONEOS) M. Jäger, Stixendorf, Austria, reports his recovery of comet P/2001

R1 (cf. IAUC 7713) with a small coma (< 20′′) and a 3′ tail in p.a. 275^{o} on CCD images taken with a 0.20-m reflector; astrometry by W. Vollmann:

2008 UT  α_{2000}  δ_{2000}  Mag.
Aug. 31.08719  7^{h}39^{m}31.^{s}16  +25^{o}12′23′′.9  16.5

New astrometry, revised orbital elements (the indicated correction to the prediction on MPC 54169 is ΔT = −0.42 day), and an ephemeris appear on MPEC 2008-R09.

2008 September 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green